OpenStudy (babe2323433):

What is the simplified form of the quantity y−squared minus y minus 12 over the quantity y−squared plus 8y plus 15?

OpenStudy (babe2323433):

@sunnystrong

OpenStudy (babe2323433):

The equation is: y^2-y-12/y^2+8y+15

OpenStudy (sunnnystrong):

So we have: \[\frac{ y^2-y-12 }{ y^2+8y+15 }\] Basically... we need to factor both quadratics \[\frac{ (y-4)(y+3) }{ (y+3)(y+5) }\] Notice that the term (y+3) cancels... so we have: \[\frac{ y-4}{ y+5 }\]

OpenStudy (babe2323433):

Oh okay

OpenStudy (babe2323433):

Time for like 5 more questions? @sunnnystrong

OpenStudy (sunnnystrong):

i will try my best :D & if not i'm sure some users would love to help lol